A 5kg block is placed on top of a 10kg block. A horizontal force of 45N is applied to the 10kg block, and the 5kg block is tied to the wall. The coefficient of kinetic friction between all moving surfaces is 0.2. Determine the (a) tension in the string and the (b) magnitude of the acceleration of the 10kg block.
1
Expert's answer
2020-10-05T10:52:53-0400

(a)


"T=\mu m_1g=0.2\cdot5\cdot9.81=9.81(N)"


(b)


"45-\mu(m_1+m_2)g-\mu m_1g=m_2a"


"a=(45-\mu(m_1+m_2)g-\mu m_1g)/m_2="


"=(45-0.2\cdot(5+10)\cdot9.81-0.2\cdot5\cdot 9.81)/10=0.576(m/s^2)"